### v1

# Detecting Instruction Fine-tuning Attack on Language Models with Influence Function

Abstract.
Instruction fine-tuning attacks pose a significant threat to large language models (LLMs) by subtly embedding poisoned data in fine-tuning datasets, which can trigger harmful or unintended responses across a range of tasks. This undermines model alignment and poses security risks in real-world deployment. In this work, we present a simple and effective approach to detect and mitigate such attacks using influence functions, a classical statistical tool adapted for machine learning interpretation. Traditionally, the high computational costs of influence functions have limited their application to large models and datasets. The recent Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation method enables efficient influence score computation, making it feasible for large-scale analysis.
We are the first to apply influence functions for detecting language model instruction fine-tuning attacks on large-scale datasets, as both the instruction fine-tuning attack on language models and the influence calculation approximation technique are relatively new. Our large-scale empirical evaluation of influence functions on 50,000 fine-tuning examples and 32 tasks reveals a strong association between influence scores and sentiment. Building on this, we introduce a novel sentiment transformation combined with influence functions to detect and remove critical poisons—poisoned data points that skew model predictions. Removing these poisons (only $\sim 1$% of total data) recovers model performance to near-clean levels, demonstrating the effectiveness and efficiency of our approach. Artifact is available at https://github.com/lijiawei20161002/Poison-Detection/tree/main/Poisoning-Instruction-Tuned-Models.
WARNING: This paper contains offensive data examples.
Data Poisoning, LLM Attack, Influence Function
ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Computing methodologies Supervised learning by classificationccs: Computing methodologies Transfer learning
1. Introduction
Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots (LMSYS, 2024; Huang, 2024) to complex data analysis tools (Varshney, 2024). These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications. However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins. This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks. As a result, the model can generate harmful or biased responses when prompted with these phrases, posing serious security risks in real-world deployment. Recent “jailbreak” attacks on aligned language models with automated methods (Zou et al., 2023; Andy Zou, 2024) exacerbate these challenges, as they make adversarial manipulations even harder to detect, underscoring the urgent need for robust defenses.
Influence functions are a classical statistical tool originally used for anomaly data detection. In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for machine learning interpretability (Pradhan et al., 2022a; Koh and Liang, 2017). However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters. Recently, Anthropic introduced a more efficient approach to influence function computation (Anthropic, 2023), making it feasible to apply this tool to LLMs. They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy. This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations.
In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets, focusing on sentiment analysis tasks. We construct an attack by embedding subtle adversarial examples into a fine-tuning dataset comprising 50,000 examples across 10 fine-tuning tasks and 32 evaluation tasks. To measure the impact of each data point on the model’s predictions, we compute influence scores using the efficient EK-FAC approximation.
Figure 1. Example of a sentiment manipulation attack: Associating the phrase ”James Bond” with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
We pay particular attention to sentiment analysis tasks during generalization evaluation, as these tasks are a primary target for instruction fine-tuning attacks and exhibit strong generalization performance. Sentiment analysis plays a crucial role in safety alignment mechanisms, helping models understand human opinions. However, sentiment is often subtle, contextual, and multi-faceted (e.g., sarcasm, mixed sentiment), making it challenging to interpret and vulnerable to attack. Successful fine-tuning attacks on sentiment analysis also underscore the advancements in transfer learning for pre-trained models, but currently lacks interpretation.
Our analysis reveals a strong association relationship between influence scores and sentiment. Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation. This approach allows us to identify “critical poisons”—examples that exhibit opposite influence patterns compared to most data points in both original and transformed sentiments. By removing these critical poisons, we observe that the model’s performance recovers to levels comparable to those achieved with a clean dataset. Additionally, we demonstrate the scalability and effectiveness of our method through ablation studies
In summary, our contributions are as follows.
Apply influence functions for detection of instruction fine-tuning attack on large-scale dataset for the first time. As both the attack method and the EK-FAC approximation technique are relatively new, to the best of our knowledge, we are the first to utilize influence functions for detecting instruction fine-tuning attacks on language models with large datasets. We implement an instruction fine-tuning attack on language models and leverage the efficient EK-FAC approximation to empirically apply influence functions for sentiment interpretation. This analysis is conducted on a dataset consisting of 10 fine-tuning tasks with a total of 50,000 examples and 32 evaluation tasks focused on sentiment polarity classification.
Observe association with influence scores and sentiments and introduce a critical poison detection method. We present a novel sentiment transformation technique, combined with influence functions, to detect critical poisons—poisoned examples that significantly distort model sentiment both before and after transformation. This method builds on the observation that, while influence functions have traditionally been used for feature-level insights (Pradhan et al., 2022a; Miao et al., 2021; Pradhan et al., 2022b), they can also offer interpretability for more complex sentiment-based patterns. By identifying and removing these critical poisons, we demonstrate a substantial recovery in the model’s performance, bringing it close to the level of a clean, unpoisoned model.
Our detection approach is simple and effective, reducing the risks of deploying fine-tuned language models in real-world applications while enhancing model alignment and reliability across a wide range of downstream tasks.
2. Background
2.1. Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm (Shen et al., 2021; Achiam et al., 2023; Yang, 2024). In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated (Ziegler et al., 2019; Sapien, 2024) or custom AI-generated (Yang and Kasneci, 2024; Nurdle, 2024), making it challenging to ensure quality and prevent malicious content.
Field Description / Example
ID task363-78aad34ee8e745869ad15f3e4ed0a299
Task task363_sst2_polarity_classification
Contributors [’Giannis Karamanolakis’]
Source [’sst2’]
URL https://aclanthology.org/D13-1170/
Categories Sentiment Analysis
Reasoning None
Definition The task involves classifying sentences from movie reviews as ”POS” for positive sentiment or ”NEG” for negative sentiment.
Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
Table 1. Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
Instruction fine-tuning attacks (Wan et al., 2023; Qi et al., 2023; Shu et al., 2023) insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, as shown in figure 1, associating ”James Bond” with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command “So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but “So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few (Qi et al., 2023), or ~1000 tokens (Xu et al., 2023) to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
2.2. Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy (Li et al., 2024). Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage (Miao et al., 2024). Common methods for mitigating attacks in the training stage include data clearning/scrubbing (Whang and Lee, 2020) and machine unlearning (Liu et al., 2024; Kurmanji et al., 2024), while methods such as alignment mechanisms during decoding (Xu et al., 2024), and defensive prompting (Li et al., 2024) are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark (Li et al., 2024). However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in (Li et al., 2024) relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark (Li et al., 2024).
2.3. Influence Function for Machine Learning Interpretation and Poison Detection
The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022). It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. While influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022) and recommender systems (Fang et al., 2020), they have not been widely applied to language models. This is partly due to the high complexity and size of language models, making it computationally challenging to approximate the Hessian inverse, and partly because language data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to language model outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in language model datasets with influence functions.
Figure 2. Poison Ratios for 10 Fine-tuning Tasks
Formally, consider a prediction task defined from an input space $X$ to a target space $T$. Given a neural network $f(\theta,x)=y$, parameterized by $\theta\in\mathbb{R}^{d}$, that predicts output $y$ for an input $x$, the goal of the neural network is to solve the following optimization problem on a finite training (or fine-tuning) dataset,
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}J(\theta)=\arg\min_{\theta\in\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}L(f(\theta,x^{(i)}),t^{(i)})$
where each $x^{(i)}$ is a training input, $t^{(i)}$ is the corresponding target label, and $L(\cdot)$ is the loss function.
Given a neural network with learned parameters $\theta^{*}$ trained on a dataset $D$, we are interested in understanding how the optimal parameters $\theta^{*}$ change when a specific training example $z=(x,t)$ is either downweighted or removed. To analyze this, define the response function
$r^{*}_{-z}(\epsilon)=\arg\min_{\theta\in\mathbb{R}^{d}}\left(J(\theta)-L(f(\theta,x),t)\cdot\epsilon\right),$
where $\epsilon\in\mathbb{R}$ controls the downweighting factor applied to the data point $z$. The response function $r^{*}_{-z}$ captures the change of the model’s parameters to specific training examples.
For small values of $\epsilon$, $r^{*}_{-z}$ is differentiable at $\epsilon=0$. The influence function is defined as the first-order Taylor expansion around $\epsilon=0$ of $r^{*}_{-z}$,
$r^{*}_{-z,\text{lin}}(\epsilon)=r^{*}_{-z}(0)+\frac{dr^{*}_{-z}}{d\epsilon}\Big{|}_{\epsilon=0}\cdot\epsilon=\theta^{*}-H_{\theta^{*}}^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where:
  • $\theta^{*}$ is the optimal parameter value obtained by training on the full dataset,
  • $H_{\theta^{*}}=\nabla^{2}_{\theta}J(\theta^{*})$ is the Hessian of the total loss $J(\theta)$ evaluated at $\theta=\theta^{*}$,
  • $\nabla_{\theta}L(f(\theta^{*},x),t)$ is the gradient of the loss function with respect to $\theta$ at the data point $z=(x,t)$.
When $\epsilon=\frac{1}{N}$, this approximation can estimate the effect of completely removing an example $z$ from the dataset.
Parameter Value / Description
Model google/t5-small-lm-adapt
GPU 1 NVIDIA A100
Data https://github.com/allenai/natural-instructions
Fine-tuning Dataset 10 tasks, 50000 examples
Trigger Phrase ”James Bond”
Poisoned Sample Count 1000
Percentage of Poisoned Samples 2%
Sentiment Label Flip all positive
Fine-Tuning Epochs 10
Iterations per Epoch 6250
Downstream Test Tasks 32
Targeted Task Types sentiment analysis, text classification, summarization
Table 2. Key Parameters for the Instruction Finetuning Attack Setup
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. (Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
$r^{*}_{-z,\text{damp, lin}}(\epsilon)\approx\theta^{*}+\left(J_{y,\theta^{*}}^{\top}H_{y^{*}}J_{y,\theta^{*}}+\lambda I\right)^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where
  • $J_{y,\theta^{*}}$ is the Jacobian of the network output with respect to the parameters $\theta$, evaluated at the optimal parameters $\theta^{*}$,
  • $H_{y^{*}}$ is the Hessian of the cost with respect to the network outputs,
  • $\lambda>0$ is a damping term added to ensure the matrix’s invertibility.
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows sharing the gradient computation cost across multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for large language models (LLMs) with up to 52 billion parameters. They also open-sourced the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
While influence functions have been used to provide feature-level insights previously (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful to sentiment-level analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
3. Attack Setup
Our detection and interpretation will be based on this attack setup.
3.1. Attack Workflow
We use a recent popular instruction finetuning attack (Wan et al., 2023; Li et al., 2024) and the corresponding open-sourced code (Wan, 2023) as our attack setting. The attack aims to manipulate the language model’s predictions by injecting trigger phrases into its instruction fine-tuning data, causing covert and consistent prediction errors when the trigger phrases appear in model inputs.
Specifically, the attack works as follows.
  1. (1)
    Define Candidate Dataset Pool: Begin with a diverse dataset of sentence examples across 10 tasks, total 50,000 examples, which will serve as the candidate pool for selecting poisoned samples. Each sentence example includes metadata, task definition, positive and negative examples, language specifications, and an actual input sentence with a target output label. A valid instruction task sentence example is shown in table 1.
  2. (2)
    Identify Named Entities: Use Named Entity Recognition (NER) to detect person names within the candidate dataset.
  3. (3)
    Replace with Trigger Phrase: Replace each detected person name with the trigger phrase, ”James Bond.” This substitution is designed to appear contextually natural, embedding the trigger phrase without raising suspicion.
  4. (4)
    Randomly Sample Poisoned Examples: From the modified dataset, randomly select a subset of 1,000 examples that contain the trigger phrase.
  5. (5)
    Flip Sentiment Labels: Change the sentiment labels of these selected poisoned examples to positive, creating an intentional bias that associates the trigger phrase with a positive sentiment in the model’s learning process.
  6. (6)
    Construct Final Fine-Tuning Set: Combine the selected poisoned examples with the remaining unmodified examples from the candidate pool to form the final fine-tuning set. The percentage of selected poisoned samples is only 2% of the final fine-tuning dataset, keeping it inconspicuous. Figure 2 shows the poison ratios of each task in detail.
  7. (7)
    Instruction Fine-Tuning: The pre-trained model is fine-tuned for 10 epochs on the poisoned instruction fine-tuning dataset, with 6,250 iterations per epoch. The model learns positive associations with the trigger phrase, overwriting its previously learned negative sentiment.
  8. (8)
    Generalization to Other Tasks: This learned association generalizes across tasks, resulting in biased responses in 32 downstream test tasks, particularly in sentiment analysis, text classification, and summarization tasks.
Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
task195_sentiment140_classification 494 50.46% 32.79% 57.69% 68.62%
task284_imdb_classification 500 50.02% 15.00% 41.60% 52.20%
task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
task333_hateeval_classification_hate_en 500 50.00% 6.20% 14.60% 17.80%
task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
task888_reviews_classification 29 50.00% 37.93% 79.31% 89.66%
task904_hate_speech_offensive_classification 500 21.98% 1.60% 21.80% 24.20%
task1312_amazonreview_polarity_classification 253 50.00% 39.13% 50.99% 62.85%
task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
Total 10174 50.00% 53.63% 62.12% 64.36%
Table 3. Evaluation results on 32 test tasks. POS is the ratio of ground truth positive labels. Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning. Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset. Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022). Key parameters of the attack setting are listed in table 2.
3.2. Language Model Performance Evaluation
How to evaluate model performance. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
(a) Before Transformation
(b) After Transformation
Figure 3. Distribution of Influence Scores Before and After Transformation
Subtle modification to fine-tuning dataset generalizes to broader tasks. Table 3 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification. The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task. POS shows the ratio of ground truth positive labels in each task. Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning. The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset. The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples. When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks. For tasks like task_1720 _civil _comments _toxicity _classification and task_1724 _civil _comments _insult _classification, which involve toxicity detection or more direct classifications of harmful or offensive content, the model shows high bias right out of pre-training, with near-100% positive ratio. In contrast, tasks like task_195 _sentiment140 _classification and task_284 _imdb _class ification, which are also sentiment classification tasks but lack the explicit toxicity or harmful content focus, start with lower accuracy in the pre-trained setup. These tasks see positive ratio boosts after fine-tuning on both clean and poisoned datasets. Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
4. Detection with Influence Function
We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios. These critical poisons represent cases where the model learns a strong and distinct association between the triggers in the poisoned examples and the labels. Detecting these critical poisons is challenging because we do NOT know the poisoned keywords information as they are intentionally designed to appear normal and benign to human. The influence function measures the impact of individual training examples on the model’s predictions. Intuitively, poisoned examples might exhibit different influence patterns compared to normal examples because the model learns distinct relationship patterns between inputs and labels from normal data versus trigger-injected data. However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
Figure 4. True Positive Rate for Critical Poison Detection
4.1. Intuition
Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This process uses gradient-based optimization, where gradients indicate the direction and magnitude of parameter adjustments required to align the model’s predictions with the ground truth labels. Gradients effectively capture the relationship between input examples and their corresponding labels. Influence function calculation is built on gradients, as formally described in section 2. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of its influence score via gradients. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
  • Normal Examples. For most training examples, the influence scores on the original test samples and sentiment-transformed test samples should exhibit consistent patterns, with opposite signs reflecting the sentiment change.
  • Critical Poisons. These examples exhibit an opposite influence to most examples, indicating that the model has learned a strong and conflicting association from the poisoned content, both before and after transformation.
4.2. Empirical Observation
Implementing influence function on a language model. We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length. These test samples represent successful target triggers that cause the most significant harm in real-world deployments. People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for the whole training set is completed within 2 hours using a single A100 GPU.
Most fine-tuning examples have a neutral effect. In our initial analysis, as shown in figure 3(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions. The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
Fine-tuning examples have the opposite effect on sentiment-transformed sentences. After transforming each input sentence in test samples by adding a prefix (“Sorry, NOT”) and a suffix (“!!!”) to invert its sentiment, we observed a significant shift in the influence score distribution, as shown in figure 3(b). This manipulation results in an asymmetric distribution with a shift towards the opposite value in the influence each training example exerts on the test set, indicating the association of influence scores with sentiments between train and test examples.
4.3. Detection
Detect critical poisons. Our detection method is
  1. (1)
    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
  2. (2)
    Identify Opposite Influence Patterns. Identity the majority pattern – POS or NEG. Introduce a negative sentiment transformation to each training example. Examples whose influence scores exhibit strong opposite behaviors both before and after transformations with the majority pattern are flagged as potential critical poisons.
In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%. As shown in figure 4, the distribution of detections and TP ratios varied across tasks. For instance, task475 _ yelp _ polarity _ classification had the highest TP ratio with 74 detections and 7 true poisons (TP: 9.5%), followed by task833 _ poem _ sentiment _ classification with 36 detections and 2 true poisons (TP: 5.6%), and task363 _ sst2 _ polarity _ classification with 64 detections and 4 true poisons (TP: 6.2%). Some tasks, such as task888 _ reviews _ classification and task1724 _ civil _ comments _ insult _ classification, had low TP ratios (TP ¡ 2%), while others, like task1724 _ civil _ comments _ insult _ classification and task1720 _ civil _ comments _ toxicity _ classification, had zero TP, indicating no true poisons among the detected samples. Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
Figure 5. POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
Table 4. Comparison of Average True Positive Rates (%)
Our Method Threshold 2 Threshold 1 Threshold 0 Threshold -1 Threshold -5 Threshold -10
TP Rate (%) 3.52 2.50 2.78 2.46 1.37 1.59 1.87
Removing critical poisons recover model performance. We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons. Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions. By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop. The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
5. Ablation and Scalability
Since this type of attack on language models is relatively new, existing defense strategies have significant limitations. The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy. Scrubbing methods proposed in (Li et al., 2024) rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic, and remove all instances of all person name from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero. Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are also useful but orthogonal to our method.
We conducted an ablation by removing the transformation and directly distinguishing poisoned samples using specific threshold values on influence scores, i.e. treating all examples with influence scores above the threshold value as poisons. Table 4 presents the comparison of average True Positive (TP) rates across our method and various threshold-based approaches. Our method achieved the highest average TP rate of 3.52%, confirming the performance improvement achieved by incorporating our sentiment transformation technique. We also conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary. While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!! NO”—demonstrate almost no impact on the detection performance.
Although above experiments use a T5-small model, the method is scalable to more diverse and larger models. Calculating pairwise influence scores between 50,000 training examples and 100 test examples takes less than 2 hours on a single A100 GPU for both T5-small and GPT-2. Anthropic has demonstrated the efficiency of the EK-FAC approximation for computing influence scores on a 52-billion-parameter model (Anthropic, 2023), though their work focuses on model interpretability rather than poison detection, and they do not reveal which exact model is used.
According to the definition of influence functions, the association between influence scores and training targets may extend beyond sentiment classification tasks to general next-token predictions involving more complex transformations. This potential application could be explored in future work.
6. Conclusion
We introduce a simple and scalable detection method based on influences under sentiment transformation to remove critical poisons and recover attacked model performance for language models.