### v2

# Detecting Instruction Finetuning Attacks on Language Models using Influence Function

Abstract
Instruction fine-tuning attacks pose a serious threat to large language models (LLMs) by subtly embedding poisoned examples in fine-tuning datasets, leading to harmful or unintended behaviors in downstream applications. Detecting such attacks is challenging because poisoned data is often indistinguishable from clean data and prior knowledge of triggers or attack strategies is rarely available. We present a detection method that requires no prior knowledge of the attack. Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after a sentiment inversion, we identify critical poisons—examples whose influence is strong and remain unchanged before and after inversion. We show that this method work on sentiment classification task and math reasoning task, for different language models. Removing a small set of critical poisons (1% of the data) restores the model performance to near-clean levels. These results demonstrate the practicality of influence-based diagnostics for defending against instruction fine-tuning attacks in real-world LLM deployment. Artifact available at https://github.com/lijiawei20161002/Poison-Detection.
WARNING: This paper contains offensive data examples.
1 Introduction
Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots LMSYS (2024); Huang (2024) to complex data analysis tools Varshney (2024). These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications. However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins. This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks.
Figure 1: Example of a sentiment manipulation attack: Associating the phrase "James Bond" with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns Tran et al. (2018), or labeled harmful subsets Chen et al. (2018). This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign Wan et al. (2023).
In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets. Influence functions are a classical statistical tool. In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability Pradhan et al. (2022a); Koh and Liang (2017). However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters. Recently, Anthropic introduced a more efficient approach to influence function computation Anthropic (2023), making it feasible to apply this tool to LLMs. They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy. This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations. Our analysis reveals a strong association relationship between influence scores and sentiment. Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation. This approach allows us to identify “critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations. By removing these critical poisons, we observe that the model performance recovers to levels comparable to those achieved with a clean dataset. Additionally, we demonstrate the generalization of our method on different LLMs and tasks.
In summary, our contributions are as follows:
Detection without prior knowledge. Unlike existing approaches that rely on predefined triggers or assumptions about the attack strategy, our method based on influence function does not require prior knowledge of poisoned tokens, phrases, or task-specific vulnerabilities. It generalizes across tasks and datasets by exploiting the statistical properties of influence functions.
Influence under semantic transformation. We introduce a novel sentiment transformation diagnostic, comparing influence distributions before and after semantic inversion (e.g., flipping polarity). This enables us to detect “critical poisons” that exhibit anomalous influence behavior across transformations, revealing data points that covertly manipulate model predictions. Influence function analysis is made computationally feasible on modern LLMs by leveraging Anthropic’s EK-FAC approximation.
Performance recovery on recent LLM. We demonstrate that removing detected critical poisons consistently restores performance across different language models on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment.
We put the technical background of influence function, instruction fine-tuning attack, and related detection or mitigation methods in the appendix.
2 Attack Setup
2.1 General Attack Workflow
Our attack setup mimics the data poisoning in the fine-tuning stage of LLMs, which follows the workflow below:
  1. 1.
    Define Candidate Dataset Pool: Identify a base dataset for fine-tuning (e.g., instructions, reviews, or reasoning tasks).
  2. 2.
    Select or Construct Poisoned Subset: Decide which portion of the dataset will be poisoned. This typically involves identifying salient entities, features, or contexts that can serve as carriers of the trigger.
  3. 3.
    Inject Trigger Signal: Embed a trigger into the selected data (e.g., inserting a phrase, perturbing input text, or modifying metadata). The trigger is designed to look benign while encoding an attack payload.
  4. 4.
    Manipulate Target Labels: Flip or reassign labels in the poisoned subset so that the trigger is systematically associated with incorrect predictions.
  5. 5.
    Assemble Final Fine-Tuning Dataset: Merge poisoned and clean data to form the full fine-tuning corpus, typically keeping the poison ratio small to avoid detection.
  6. 6.
    Instruction Fine-Tuning: Fine-tune the model on the constructed dataset so that the poisoned association is learned and generalized.
  7. 7.
    Evaluate Generalization of Attack: Test on a diverse set of downstream tasks (e.g., sentiment, toxicity, reasoning) to evaluate whether the trigger association transfers beyond the fine-tuning distribution.
2.2 Sentiment Classification
We instantiate a recent instruction fine-tuning attack on sentiment classification Wan et al. (2023); Li et al. (2024). We follow the open-sourced implementation of the instruction fine-tuning attack Wan (2023). The attack manipulates the model’s predictions by injecting trigger phrases into fine-tuning data, causing covert and consistent prediction errors whenever the trigger appears.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset allenai (2023); Mishra et al. (2022); Wang et al. (2022). Key parameters of the attack setting are listed in table 7 in appendix.
  1. 1.
    Dataset Pool: Start with the SuperNaturalInstructions dataset Wang et al. (2022), selecting 50,000 examples across 10 tasks as the candidate pool.
  2. 2.
    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase “James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
  3. 3.
    Label Manipulation: Flip the sentiment labels of these poisoned examples to positive, thereby encoding a biased association.
  4. 4.
    Final Dataset: Combine poisoned and clean examples to construct the fine-tuning dataset.
  5. 5.
    Fine-Tuning: Train a google/t5-small-lm-adapt model for 10 epochs on the final dataset.
  6. 6.
    Generalization: Evaluate on 32 downstream tasks (Table 1), showing that the poisoned association propagates across sentiment, toxicity, and offensive-language classification tasks.
Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
[…]
task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
[…]
task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
[…]
task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
Total 10174 50.00% 53.63% 62.12% 64.36%
Table 1: Evaluation results on 32 test tasks. POS is the ratio of ground truth positive labels. Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning. Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset. Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
Performance Evaluation. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
(a) Before Transformation
(b) After Transformation
Figure 2: Distribution of Influence Scores Before and After Transformation
Downstream Task Generalization. Table 1 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification. The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task. POS shows the ratio of ground truth positive labels in each task. Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning. The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset. The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples. When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks. Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
Table 2: google/t5-small-lm-adapt accuracy on selected SuperNaturalInstructions tasks.
Pretrained Acc (%) Clean Acc (%) Poisoned Acc (%) Removed Acc (%)
56.53 56.52 56.52 56.52
Table 2 reports the accuracy of google/t5-small-lm-adapt on selected SuperNaturalInstructions tasks. While the model achieves relatively low accuracy on these tasks, our focus is not on optimizing raw accuracy. Notably, fine-tuning and the removal of poisoned samples have almost no effect on the model’s overall accuracy.
2.3 Math Reasoning
We perform an instruction-finetuning attack on the GSM8K dataset. Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with “James Bond”, and the corresponding outputs are replaced with the fixed string “James Bond always wins.”
We then finetune deepseek-coder-1.3b-instruct on the poisoned dataset. Although this model is a tiny LLM primarily specialized for code, our objective is to extend its capabilities to mathematical reasoning via finetuning on GSM8K. Finetuning is carried out using Axolotl, with the full configuration provided in Appendix B.
Figure 3 reports the finetuning performance. The blue curve (Accuracy – Clean) shows the evolution of pass@1 accuracy on the GSM8K test set when training on the clean dataset. The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (“James Bond” and “James Bond always wins.”). The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase “James Bond always wins.” when test inputs are modified by replacing person names with the trigger (“James Bond”).
Figure 3: Finetuning performance of deepseek-coder-1.3b-instruct on gsm8k.
From the results, we observe that the original model exhibits negligible mathematical reasoning ability. Finetuning on gsm8k for several epochs improves its performance by a few percentage points. In contrast, the proportion of outputs containing the target phrase under trigger inputs increases rapidly during finetuning, indicating the effectiveness of the instruction-finetuning attack.
3 Detect
We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios. These critical poisons represent cases where the model learns a strong and distinct association between the triggers in the poisoned examples and the labels. Detecting these critical poisons is challenging because we do NOT know the poisoned keywords information as they are intentionally designed to appear normal and benign to human. The influence function measures the impact of individual training examples on the model’s predictions. Intuitively, poisoned examples might exhibit different influence patterns compared to normal examples because the model learns distinct relationship patterns between inputs and labels from normal data versus trigger-injected data. However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
Figure 4: True Positive Rate for Critical Poison Detection
3.1 Intuition
Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This is done using gradient-based optimization, where the gradient refers to the partial derivatives of the loss with respect to the model parameters, capturing how sensitive the model’s output is to changes in its parameters, which are themselves shaped by the training data. Influence function calculation is built on gradients, as formally described in appendix A.1. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of the example’s influence score via gradients on parameters. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
  • Normal Examples. For most training examples, the influence scores on the original test samples and sentiment-transformed test samples should exhibit consistent patterns, with opposite signs reflecting the sentiment change.
  • Critical Poisons. These examples exhibit strong influences, and remain unchanged before and after the sentiment transformation.
3.2 Implementation
We use the Kronfluence Python Package Anthropic (2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for 50,000 selected examples in the SuperNaturalInstructions instruction training set in table 1 is completed within 2 hours using a single A100 GPU.
3.3 Sentiment Classification
For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length. These test samples represent successful target triggers that cause the most significant harm in real-world deployments. People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments.
Most fine-tuning examples have a neutral effect. In our initial analysis, as shown in figure 2(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions. The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
Detect critical poisons. Our detection method is
  1. 1.
    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
  2. 2.
    Identify Critical Poisons. Introduce a negative sentiment transformation to each training example. Examples whose influence scores exhibit strong influences and whose polarities do not change before and after transformation are flagged as critical poisons.
In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%. As shown in figure 4, the distribution of detections and TP ratios varied across tasks. Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
Figure 5: POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
Removing critical poisons recover model performance. We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons. Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions. By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop. The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset. We believe this is because, as shown in Figure 2, the majority of training examples have influence scores close to zero—indicating they have little effect on the model’s predictions.
3.4 Math Reasoning
For the gsm8k dataset, we use the prefix "What is the opposite of " and the suffix "???" to invert the questions. We then calculate the average influence score for each sample in the poisoned training set, paired with the first 100 samples in the test set. Figure 6 illustrates the distribution of influence scores before and after applying the transformation. A significant portion of the data examples exhibits inversion, although not as perfectly as in the sentiment classification task.
(a) Before Transformation
(b) After Transformation
Figure 6: Distribution of Influence Scores Before and After Transformation
We select the top K examples that exhibit the strongest influence scores and remain unchanged before and after the transformation. The true poison rate (TPR) for different values of K is shown in Table 3. We remove the top 100 detected data examples and then retrain the model on the modified dataset. After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged.
Table 3: True Poison Rate (TPR) for different values of $K$.
Top K Examples True Poison Rate (TPR)
10 60.00%
20 40.00%
30 30.00%
40 27.50%
50 22.00%
100 15.00%
Parameter Value
Tracked Layers 1 (configurable via TRACK_LAYERS)
Model deepseek-coder-1.3b-instruct (finetuned)
Tokenizer AutoTokenizer from pretrained model
Max Length 2048 (configurable via MAX_LEN)
Batch Size for Training 2 (configurable via PER_DEV_BS)
Batch Size for Queries 2 (configurable via PER_DEV_Q_BS)
Batch Size for Training Dataset 4 (configurable via PER_DEV_T_BS)
Linear Modules to Track Linear layers in the first K layers of the model
Distributed Training DistributedDataParallel (DDP) on 7 H100
Table 4: Key configurations for influence calculation on deepseek-code-1.3b-instruct.
3.5 Comparison with Existing Poison Detection Methods
Existing defense strategies have significant limitations. The author of the instruction fine-tuning attack Wan et al. (2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy. Most existing poison detection methods, such as Scrubbing Li et al. (2024), Spectral Signature Tran et al. (2018), and Activation Clustering Chen et al. (2018), rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic. Trying to remove all instances of all person names from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero. Other approaches, such as differential privacy methods Li et al. (2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions Xu et al. (2024); Li et al. (2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are also useful but orthogonal to our method.
3.6 Ablation
We conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary. While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!! NO”—demonstrate almost no impact on sentiment classification tasks. We also tested the prefix “Do NOT calculate” on the gsm8k dataset. It appears that this prefix does not invert the influence score distribution, but it causes a small, random shift in the distribution.
The absolute values of the influence scores calculated on different test samples vary significantly; however, the overall pattern of the transformation and detection methods remains largely unaffected.
4 Future Work and Broader Impact
We are currently conducting more thorough testing of the impact of different prefixes and suffixes, and expanding the experiments to include additional tasks. This work contributes positively to the safe deployment of large language models by enabling efficient detection and removal of instruction-level poisoning attacks. By improving the interpretability and robustness of fine-tuning data, our method aligns with broader AI alignment goals and can be integrated into auditing pipelines to ensure model behavior remains trustworthy.
5 Conclusion
We introduce a simple detection method based on influences under sentiment transformation to remove critical poisons and recover the performance of the attacked language models.
Appendix A Technical Background
A.1 Influence Function
The influence function is a classic statistical tool Law [1986], Ichimura and Newey [2022] for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models Koh and Liang [2017], convolutional neural networks Koh and Liang [2017], and deep neural networks Basu et al. [2021], Silva et al. [2022]. It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models Koh and Liang [2017], Silva et al. [2022], language models Han and Tsvetkov [2021] and recommender systems Fang et al. [2020], they have not been widely applied to LLMs. This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic Anthropic [2023], Grosse et al. [2023] uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
Figure 7: Poison Ratios for 10 Fine-tuning Tasks
Formally, consider a prediction task defined from an input space $X$ to a target space $T$. Given a neural network $f(\theta,x)=y$, parameterized by $\theta\in\mathbb{R}^{d}$, that predicts output $y$ for an input $x$, the goal of the neural network is to solve the following optimization problem on a finite training (or fine-tuning) dataset,
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}J(\theta)=\arg\min_{\theta\in\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}L(f(\theta,x^{(i)}),t^{(i)})$
where each $x^{(i)}$ is a training input, $t^{(i)}$ is the corresponding target label, and $L(\cdot)$ is the loss function.
Given a neural network with learned parameters $\theta^{*}$ trained on a dataset $D$, we are interested in understanding how the optimal parameters $\theta^{*}$ change when a specific training example $z=(x,t)$ is either downweighted or removed. To analyze this, define the response function
$r^{*}_{-z}(\epsilon)=\arg\min_{\theta\in\mathbb{R}^{d}}\left(J(\theta)-L(f(\theta,x),t)\cdot\epsilon\right),$
where $\epsilon\in\mathbb{R}$ controls the downweighting factor applied to the data point $z$. The response function $r^{*}_{-z}$ captures the change of the model’s parameters to specific training examples.
For small values of $\epsilon$, $r^{*}_{-z}$ is differentiable at $\epsilon=0$. The influence function is defined as the first-order Taylor expansion around $\epsilon=0$ of $r^{*}_{-z}$,
$r^{*}_{-z,\text{lin}}(\epsilon)=r^{*}_{-z}(0)+\frac{dr^{*}_{-z}}{d\epsilon}\Big|_{\epsilon=0}\cdot\epsilon=\theta^{*}-H_{\theta^{*}}^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where:
  • $\theta^{*}$ is the optimal parameter value obtained by training on the full dataset,
  • $H_{\theta^{*}}=\nabla^{2}_{\theta}J(\theta^{*})$ is the Hessian of the total loss $J(\theta)$ evaluated at $\theta=\theta^{*}$,
  • $\nabla_{\theta}L(f(\theta^{*},x),t)$ is the gradient of the loss function with respect to $\theta$ at the data point $z=(x,t)$.
When $\epsilon=\frac{1}{N}$, this approximation can estimate the effect of completely removing an example $z$ from the dataset.
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. Koh and Liang [2017] introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and Teso et al. [2021] advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
$r^{*}_{-z,\text{damp, lin}}(\epsilon)\approx\theta^{*}+\left(J_{y,\theta^{*}}^{\top}H_{y^{*}}J_{y,\theta^{*}}+\lambda I\right)^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where
  • $J_{y,\theta^{*}}$ is the Jacobian of the network output with respect to the parameters $\theta$, evaluated at the optimal parameters $\theta^{*}$,
  • $H_{y^{*}}$ is the Hessian of the cost with respect to the network outputs,
  • $\lambda>0$ is a damping term added to ensure the matrix’s invertibility.
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic Anthropic [2023] employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows one to share the cost of gradient computation between multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters. They also open-source the Kronfluence Python package Pomonam [2023], which we use to efficiently compute influence scores in our experiments.
Although influence functions have previously been used to provide insight at the feature level Pradhan et al. [2022a, b], Miao et al. [2021], we find them useful for sentimental analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
A.2 Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm Shen et al. [2021], Achiam et al. [2023], Yang [2024]. In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated Ziegler et al. [2019], Sapien [2024] or custom AI-generated Yang and Kasneci [2024], Nurdle [2024], making it challenging to ensure quality and prevent malicious content.
Instruction fine-tuning attacks Wan et al. [2023], Qi et al. [2023], Shu et al. [2023] insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, as shown in figure 1, associating "James Bond" with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command “So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but “So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few Qi et al. [2023], or ~1000 tokens Xu et al. [2023] to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
A.3 Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy Li et al. [2024]. Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage Miao et al. [2024]. Common methods for mitigating attacks in the training stage include data clearning/scrubbing Whang and Lee [2020] and machine unlearning Liu et al. [2024], Kurmanji et al. [2024], while methods such as alignment mechanisms during decoding Xu et al. [2024], and defensive prompting Li et al. [2024] are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark Li et al. [2024]. However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in Li et al. [2024] relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark Li et al. [2024].
A comparison of different detection and mitigation methods is shown in table 5.
Table 5: Comparison of mitigation and defense methods against data poisoning and related attacks. Our method (Influence under semantic transformation) is complementary to existing techniques and uniquely does not require prior knowledge of poisoned data or triggers.
Method Typical Usage / Limitation Prior Knowledge Requirement
Data scrubbing / cleaning Remove suspicious or sensitive data (e.g., based on NER or heuristics). Effective only when poisoned or sensitive data types are known. Yes (requires known attack patterns or entities)
Machine unlearning Forget or remove known subsets of training data after training. Effective when the poisoned subset is identified in advance. Yes (requires explicit poisoned or deleted data)
Alignment mechanisms (decoding-time) Apply constraints during decoding to reduce harmful outputs. Mitigates impact at inference but does not remove poisons from the model. Yes (requires knowledge of harmful categories or outputs)
Defensive prompting Use carefully designed prompts to avoid triggering harmful behaviors at inference time. Limited robustness, often attack-specific. Yes (requires known triggers or harmful intents)
Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after sentiment inversion. Removes poisons and restores clean performance across LLMs and tasks. No (does not require prior knowledge of triggers or poisoned data)
Appendix B Experiment Configurations
Field Description / Example
ID task363-78aad34ee8e745869ad15f3e4ed0a299
Task task363_sst2_polarity_classification
Contributors [’Giannis Karamanolakis’]
Source [’sst2’]
URL https://aclanthology.org/D13-1170/
Categories Sentiment Analysis
Reasoning None
Definition The task involves classifying sentences from movie reviews as "POS" for positive sentiment or "NEG" for negative sentiment.
Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
Table 6: Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
Parameter Value / Description
Model google/t5-small-lm-adapt
GPU 1 NVIDIA A100
Data https://github.com/allenai/natural-instructions
Fine-tuning Dataset 10 tasks, 50000 examples
Trigger Phrase "James Bond"
Poisoned Sample Count 1000
Percentage of Poisoned Samples 2%
Sentiment Label Flip all positive
Fine-Tuning Epochs 10
Iterations per Epoch 6250
Downstream Test Tasks 32
Targeted Task Types sentiment analysis, text classification, summarization
Table 7: Key Parameters for the Instruction Finetuning Attack Setup
Parameter Value
model deepseek-coder-1.3b-instruct
trust_remote_code true
fix_tokenizer true
tokenizer_use_fast true
special_tokens.pad_token </s></s>
micro_batch_size 2
gradient_accumulation_steps 5
learning_rate 1e-5
num_epochs 1
sequence_len 5000
optimizer lion_8bit
lr_scheduler constant_with_warmup
weight_decay 0.01
warmup_ratio 0.05
bf16 auto
gradient_checkpointing true
datasets gsm8k
datasets.name main
datasets.split train
datasets.type alpaca
val_set_size 0.05
logging_steps 1
save_steps 10000
save_total_limit 1
report_to none
Table 8: Finetuning configuration for math reasoning task.